Demand and supply functions of a commodity are given respectively as quantity demanded 40 at 400, quantity supply 120 naira per 24 unit. Determine the equilibrium of quantity bought

1
Expert's answer
2022-01-16T17:43:35-0500

"Q_d=(400,0), (0,40)"

"Demand function= y=x-mx"

"m=\frac {(400-0)} {0-40}=-10"

x= 40 ( Y intercept)

"Q_d=40-10P"


"Q_s=(0,120), (24,0)"

"Supply function= y=x+mx"

"m=\frac {(24-0)} {0-120}= 0.2"

"Q_s=24+0.2P"

At Equilibrium, "Q_d=Q_s"

40-10P=24+0.2P

10.2P=16

P=1.6

Equilibrium Quantity

"40- (10\times1.6)=24"